President of the PTA at a local elementary school is concerned that Elemenatry students are spending too much time ONLINE. The President does a survey of middle and highschool students too, to compare number of hours spent online perday. They measure how many hours per day in total are spent online for social media, school work, entertainment or other reasons. The results of the chi Square test of association are shown below. Chi-Square Test for Association: C14, Worksheet columns Rows: C14 Columns: Worksheet columns 9 or More Hrs Five-Nine Two-Five Two orLess All Elementary 25 33 44 12 114 26.36 35.26 42.11 10.27 0.07020 0.14501 0.08500 0.29132 Middle 23 39 38 10 110 25.44 34.02 40.63 9.91 0.23319 0.72773 0.17032 0.00082 HighSchool 29 31 41 8 109 25.20 33.71 40.26 9.82 0.57165 0.21859 0.01355 0.33725 All 77 103 123 30 333 Cell Contents Count Expected count Contribution to Chi-square Chi-Square Test Chi-Square DF P-Value Pearson 2.865 X 0.826 Likelihood Ratio 2.834 X 0.829

What is the Chi Square degrees of freedom?

9

6

8

12
Answer #1

Degrees of freedom = (Number of rows - 1)*(Number of columns - 1) = (4 - 1) * (3 - 1) = 6

Option B is correct.